A bus is moving at 10ms^-1 towards a traffic light.when the driver sees the light changed to red,the bus is at a distance of 20m.the reaction time of driver us 0.5s and the deceleration is 4.5ms^-2.with the brake fully applied,how far is the bus when stops?
1
Expert's answer
2014-07-01T05:15:49-0400
Time of driver's reaction is 0,5 s. Hence, the distance tothe light when the bus will actually start to brake is 20-10*0.5=15m
Also we need to find a time to stop thebus. v=a*t. 10=4.5*t,t=20/9s
Now, we can calculate the distance travelled after theapplication of brake: S=v*t-a*t^2/2=10*20/9-(4.5/2)*(20/9)^2=11(1/9)m

So, the distance is: 15-11(1/9)=3(8/9)m=3.889 m